OpenStudy (anonymous):

\[MULTIPLY (2\sqrt{3}-1)(3\sqrt{3}-2)\]

OpenStudy (anonymous):

This is easy you can do it

OpenStudy (anonymous):

if i could i wouldnt have posted it

OpenStudy (anonymous):

help..

OpenStudy (anonymous):

7.87564

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[(2\sqrt{3}-1)(3\sqrt{3}-2)\] \[2\sqrt{3}\times 3\sqrt{3}-2\times 2\sqrt{3}-3\sqrt{3}+2\] \[6\times 3-7\sqrt{3}+2\] \[18-7\sqrt{3}+2\] \[20-7\sqrt{3}\]